### v2

# RAG-QA Arena: Evaluating Domain Robustness for Long-Form Retrieval-Augmented Question Answering

Abstract
Question answering based on retrieval-augmented generation (RAG-QA) is an important research topic in NLP and has a wide range of real-world applications. However, most existing datasets for this task are either constructed using a single source corpus or consist of short extractive answers, which fall short of evaluating large language model (LLM) based RAG-QA systems on cross-domain generalization. To address these limitations, we create Long-form RobustQA (Lfrqa), a new dataset comprising human-written long-form answers that integrate short extractive answers from multiple documents into a single, coherent narrative, covering 26K queries and large corpora across seven different domains. We further propose Rag-qa Arena by directly comparing model-generated answers against Lfrqa’s answers using LLMs as evaluators. We show via extensive experiments that Rag-qa Arena evaluation and human judgments on answer quality are highly correlated. Moreover, only 41.3% of the most competitive LLM’s answers are preferred to Lfrqa’s answers, demonstrating Rag-qa Arena as a challenging evaluation platform for future research. [^1] Code: https://github.com/awslabs/rag-qa-arena
RAG-QA Arena: Evaluating Domain Robustness for Long-Form Retrieval-Augmented Question Answering
Rujun Han2†   Yuhao Zhang3†   Peng Qi4†   Yumo Xu1   Jenyuan Wang1 Lan Liu1   William Yang Wang 5†  Bonan Min1  Vittorio Castelli1 1AWS AI Labs  2Google  3Samaya.ai  4Orby.ai  5University of California, Santa Barbara rujunh@google.com; yuhao@samaya.ai; peng@orby.ai; william@cs.ucsb.edu {liuall,yumomxu,jenwan,bonanmin,vittorca}@amazon.com
footnotetext: Work done at AWS AI Labs.
1 Introduction
Figure 1: Lfrqa annotation example. There are three documents (some text removed for brevity) relevant to the query. We instruct annotators to combine Robustqa’s answers into a coherent long-form answer with added text if necessary. Citations [1], [2] and [3] indicate the supporting documents of each sentence.
Dataset Answers grounded Long-form Multiple Coherent Multiple Human # Test
Name in corpus answers documents answers domains annotated queries
Lfrqa 16.1K
RobustqaHan et al. (2023) 16.1K
NQ Kwiatkowski et al. (2019) 03.6K
MultiHop-RAG Tang and Yang (2024) 02.5K
ASQA Stelmakh et al. (2022) 01.0K
LongFact Wei et al. (2024) 02.3K
ELI5 Fan et al. (2019) 25.0K
Table 1: Comparison of datasets. Lfrqa distinguishes from previous work by uniquely encompassing seven features: 1) RAG-QA dataset with answers annotated based on underlying corpus; 2) Long-form answers of paragraph length; 3) Multiple documents that provide different facts/views; 4) Coherent answers that handle conflicting information; 5) Multiple-domain corpus to benchmark domain robustness; 6) Human annotated high-quality answers; 7) Large-scale evaluation set.
Traditional reading comprehension task is constrained to fixed contexts Rajpurkar et al. (2016); Kočiský et al. (2018); Huang et al. (2019). It is inadequate at addressing real-world questions, where no context is readily provided for a system to find answers. Such open-ended questions require a system to identify answers in an enormous knowledge base (e.g., Wikipedia) that is computationally prohibitive to feed into question answering systems such as large language models (LLMs). Retrieval-augmented generative question answering (RAG-QA) becomes an effective tool to filter out massive amounts of noise and select only a few highly relevant passages for LLM-based QA models.
The wide applications of RAG-QA Gao et al. (2023) necessitate the evaluation of systems’ out-of-domain (OOD) performances, because a real-world system often confronts new data unseen during training. Existing popular benchmark datasets such as Natural Questions (NQ) Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017) consist solely of Wikipedia or Web documents, which fall short at measuring OOD performances.
Robustqa Han et al. (2023) was the first dataset created to benchmark cross-domain robustness for RAG-QA. However, as illustrated by the yellow highlights in Figure 1, Robustqa follows NQ’s annotation format with short answer spans extracted from the documents. Such data format is not the most suitable reference answer to evaluate the current leading LLMs that typically generate long-form responses with multiple pieces of information combined in one coherent narrative Brown et al. (2020); Ouyang et al. (2022); OpenAI (2024). Consequently, token overlap metrics used in the extractive QA era Karpukhin et al. (2020); Izacard et al. (2024) will penalize unfairly on the additional supporting tokens generated by LLMs, resulting in extremely low overlap scores. As an example, Fig. 1’s extractive answers have poor Exact-Match or $F_{1}$ scores with the final long-form answer. To create a long reference answer, one could simply concatenate these short answers, but the synthesized answers are either incoherent or ill-formatted, as illustrated by examples in Sec. 3.1.
To address these drawbacks, we propose long-form RobustQA (Lfrqa) that integrates multiple short extractive answers into a coherent long-form answer. Figure 1 shows an annotation where three extractive answers are combined by annotators to create a comprehensive answer. Table 1 summarizes seven features in Lfrqa  that make it uniquely beneficial for RAG-QA evaluations. ASQA Stelmakh et al. (2022) and ELI5 Fan et al. (2019) are the most similar datasets to Lfrqa. However, they are either not directly annotated against the underlying corpus (thus, not RAG-QA), or rely on single-domain data, which is insufficient to benchmark systems’ cross-domain performances.
With Lfrqa annotations, we propose Rag-qa Arena that leverages model-based evaluators to directly compare LLMs’ answers with Lfrqa without the necessity to examine long and potentially noisy retrieved passages. By demonstrating the high correlation with human judges following the same instruction and rubrics, we show that Rag-qa Arena is an efficient and accurate framework to benchmark the RAG-QA system’s cross-domain performances. In this work, we primarily focus on the LLMs used for the QA component, but Rag-qa Arena can be easily extended to study retrieval’s impact on answer generation quality.
We summarize our contributions: 1) We present Lfrqa, the first high-quality and large-scale multi-domain human annotations with coherent long-form answers for RAG-QA. 2) We propose an efficient model-based evaluation framework, Rag-qa Arena that enables users to directly compare LLMs’ answers with ground-truth answers in Lfrqa. 3) We build a dashboard incorporating a wide range of leading LLMs and conduct in-depth analysis to show that Lfrqa’s answers are preferred significantly more to the best LLMs with long context. Therefore, we believe Rag-qa Arena will serve as a challenging and robust evaluation benchmark for future RAG-QA research.
Figure 2: Lfrqa v.s. Robustqa. Citations are removed in Lfrqa’s answers, and a few answer spans are removed for clarity. Green and orange texts represent positive and negative opinions, respectively.
2 RAG-QA Task Formulation
We briefly introduce the RAG-QA task in this section. Passage retrieval is the first step of a RAG-QA pipeline. Following the passage retrieval set-up in DPR Karpukhin et al. (2020) and Robustqa Han et al. (2023), we denote a collection of documents as $\mathcal{D}$. We split each document $d^{i}\in\mathcal{D}$ with a fixed length $N$ tokens and obtain a collection of $M$ ($\geq|\mathcal{D}|$) passages denoted as $\mathcal{C}=\{p_{1},p_{2},...p_{m},...p_{M}\}$, where $p_{m}$ is a passage. Given a question $q$, the passage retrieval task is to select $K$ most relevant passages for $q$ with a retriever $\mathcal{R}$ from $\mathcal{C}$. Formally, $\mathcal{R}(q,\mathcal{C})\rightarrow\mathcal{C}_{q}$.
Upon receiving top $K$ passages or $\mathcal{C}_{q}$, a QA model reads them as context to generate an answer for the query. Unlike the extractive QA setting in Robustqa and NQ, we adopt generative QA as it is most compatible with the generative nature of the leading LLMs with the flexibility to produce free-form answers. The answer generation task can be modeled as $\Sigma_{1}^{T}\mathcal{P}_{q}(w_{t}|w_{0:t-1};\mathcal{C}_{q})$, where $\mathcal{P}$ is an LLM. We focus on the variations of $\mathcal{P}$ and fix $\mathcal{R}$ in this work.
In real-world applications, we deploy RAG-QA systems into various domains such as healthcare, finance, and technology whose corpus and query types may not be well covered in a trained retriever and LLM. Lewis et al. (2021) and Han et al. (2023) show significant performance gaps between in-domain and out-of-domain data for RAG-QA systems, further verifying domain adaptation problems. Therefore, it is crucial to gauge the domain robustness of RAG-QA based on LLMs, and Lfrqa helps achieve this evaluation goal.
3 Data Creation
Test Set Robustqa Lfrqa
Domain Source Label $|Q|$ $|D|$ $|P|$ $A/Q$ $W/A$ $A/Q$ $W/A$
Biomedical BioASQ [BI] 1,956 15,559,026 37,406,880 2.6 2.4 1.0 30.0
Finance FiQA [FI] 3,612 57,638 105,777 3.0 9.4 1.0 69.1
Lifestyle LoTTE [LI] 2,208 119,461 241,780 5.7 8.7 1.0 99.5
Recreation LoTTE [RE] 2,094 166,975 315,203 3.2 7.2 1.0 60.3
Technology LoTTE [TE] 2,111 638,509 1,252,402 6.0 8.7 1.0 99.7
Science LoTTE [SC] 1,423 1,694,164 3,063,916 5.3 7.8 1.0 92.0
Writing LoTTE [WR] 2,695 199,994 347,322 6.2 6.6 1.0 88.0
Table 2: Data (test set) summary: Lfrqa v.s. Robustqa. $|Q|$, $|D|$, $|P|$, $A/Q$, and $W/A$ represent numbers of questions, documents, passages, answers per question, and words per answer, respectively. Each passage consist of 100 words at most. Lfrqa  has only one answer per query as we integrate multiple answers from Robustqa, which results in more words in (long-form) answers. Dev set statistics can be found in Appendix Table 7.
Lfrqa consists of two types of QA samples: 1) new annotations in Finance ([FI]), Lifestyle ([LI]), Recreation ([RE]), Technology ([TE]), Science ([SI]), and Writing ([WR]) domains; 2) adapted long-form BioASQ ([BI]). We describe the details of both QA samples in the following sections.
3.1 Annotated Data
Following Robustqa Han et al. (2023), Lfrqa’s new annotations are also based on the LoTTE and FiQA queries and corpus. LoTTE was proposed in the ColBERTv2 paper Santhanam et al. (2022) and consists of information retrieval (IR) datasets across five domains: lifestyle, recreation, technology, writing, and science, each can have relevant answers coming from either web search or on-line forum. FiQA Maia et al. (2018) proposes a task, “Opinion-based QA over financial data” that answers finance-related questions from financial corpora such as microblogs, reports, and news. It is important to note that both FiQA and LoTTE are IR datasets with answers as long documents, which may include a large amount of irrelevant information to the query.
As IR datasets, both FiQA and LoTTE could only provide relevant documents to users, as there are no precise answer annotations. Robustqa addresses this short-coming by extracting short answer spans from the long documents in the similar format of NQ, which serves as a high-quality benchmark for extractive RAG-QA. Figure 1 shows an example where the yellow highlights in Documents 1-3 are the extracted answers to the question.
Limitations of extractive RAG-QA.  In the era of LLMs, models’ responses to user queries are often long and comprehensive OpenAI (2024); Anthropic (2024); MetaAI (2024); Jiang et al. (2024), which the short, extractive reference answers in Robustqa are no longer the most compatible format to evaluate against. First, in Robustqa, annotators are limited to only taking 3 answer spans per relevant document, each with no more than 16 words. This process could result in a loss of useful information to help answer the query. Second, to reconcile multiple extractive references for model evaluation, prior work in extractive RAG-QA Karpukhin et al. (2020); Han et al. (2023); Izacard et al. (2024) adopt the maximum of token overlaps between a model prediction and a list of references to compute EM or $F_{1}$ metrics, which penalizes unfairly the long-form responses from modern LLMs. Finally, if we naively concatenate or list all short answer spans as shown by examples in Figure 2, the combined answers are often too ill-formatted or incoherent as ground-truth answers.
Lfrqa addresses all of these drawbacks by instructing the annotators to integrate all short answers in Robustqa into a coherent long-form answer. Below, we show a summary of our annotation instruction and quality control mechanism.
Annotation instruction.  As Fig. 1 shows, a query, all relevant documents, and original short answers (highlighted in the documents) are presented to annotators on a single annotation page. Annotators need to combine all highlighted answers into a single complete and coherent answer. All highlighted answers MUST be included; otherwise, the annotation is considered as a failure. Annotators are encouraged to include more information in the documents if it helps to answer the queries. To ensure annotators faithfully use the document information, we request annotators to provide citations after each answer sentence. For example, the first sentence in Fig. 1 is composed using information from Documents 2 and 3. Annotators should add "[2, 3]" after that sentence. We use these citations primarily for data quality control and remove them during the answer evaluation. The actual annotation UI can be found in Appendix Fig. 5.
Quality control.  The data annotations are performed by contracted data professionals. We also have a dedicated team of data linguists to validate the annotation quality. Specifically, our data linguists randomly audit 10% of each batch of the annotations, and if the valid answer ratio is $<90\%$, we send the batch back to the annotators for re-work. The process iterates until the valid answer ratio exceeds $90\%$. Here is a list of failure cases: 1. Incompleteness:  Answers do not include all highlighted answers, or there is clear relevant information in the documents, but not included in the answer. 2. Redundancy:  Clear irrelevant information is included in the answer. 3. Incoherence:  Answers are not coherent or not written in natural English. 4. Citation Error:  Wrong/missing citations, which indicate annotators do not use correct information from the right documents.
3.2 Adapted Data
For the biomedical domain, Lfrqa leverages the same set of test queries as in Robustqa, but uses the complete rather than span answers in the BioASQ dataset. The original BioASQ annotations provide two types of answer formats: 1) exact answer, which is the short extractive answers used in Robustqa; 2) ideal answer, which is a long-form abstractive answer to be consistent with other datasets in this work. We did not perform further annotations. We notice in Table 2 that BioASQ’s answers are shorter compared with other datasets. This is due to its dominant amount of factoid queries, which do not require elaborated explanations as in other datasets with more open-ended reasoning questions Han et al. (2023).
We drop SearchQA in Robustqa as this dataset only has short-form extractive answers, and its documents contain a significant amount of text omission (“…”) that prevents us from re-constructing long-form answers.
3.3 Data Statistics and Analysis
Table 2 summarizes the statistics for the test set, which consists of 16K queries across 7 domains. We filter out queries with more than 80 ground-truth documents, resulting in 73 fewer queries compared with Robustqa. Since Lfrqa combines multiple short answers, the answer per query ratio (A/Q) is always 1, and the word per answer ratio (W/A) is substantially higher compared with Robustqa. We also annotate a dev set with 10K queries for future model development purposes, and the statistics can be found in Appendix Table 7. We conduct further analysis below to demonstrate the unique contributions of Lfrqa in Table 1.
Answers over multiple documents.  Figure 3 illustrates the distribution of number of documents used by Lfrqa’s answers. Specifically, Figure 3(a) shows that around 65% of the answers use $\geq$ 2 documents’ information. 4.9% of the answers consist of information from 10 or more documents (maximum = 80). In Figure 3(b), we divide long-form answers into sentences and show the distribution of the number of documents used per answer sentence. Nearly 22% of the answer sentences combine information from multiple documents. Both show that Lfrqa’s answers effectively combine information across multiple ground-truth documents. This makes Lfrqa  challenging for RAG-QA, as it requires identification and aggregation of information across sources.
Coherent answers.  Lfrqa’s answers further organize facts and views across multiple documents in a coherent paragraph. Answers with multiple views are common in the original Robustqa’s answers. Conducting a string match of both “yes” and “no” as a leading word in an answer list, we found more than 200 examples with such conflicting information. This does not account for more subtle cases where answers semantically contradict each other. In Figure 2, we show 2 examples with conflicting views. Robustqa’s annotations simply list them (separated by new-lines), whereas Lfrqa’s answers organize them as coherent narratives, with conflicting information reconciled in helpful context.
Fluency.  Robustqa’s answers are extracted from documents and often cut off unnaturally to satisfy the limit of 16 words. In Figure 2, Robustqa’s answers, such as “many ways to compromise your identity” and “mostly anecdotal evidence here suggest no” are incomprehensible without further context, whereas Lfrqa’s answers are all well written in complete sentences.
All of these features show that Lfrqa provides both challenging and high-quality annotations for evaluating RAG-QA systems.
(a) # of documents used per answer.
(b) # of documents used per answer sentence.
Figure 3: Distribution of number (#) of documents used in Lfrqa’s answers. All numbers are %.
4 RAG-QA Arena
In this section, we propose our evaluation framework Rag-qa Arena. Inspired by the pairwise human preference evaluation framework such as Chatbot Arena Chiang et al. (2024), we calculate win-rate and win+tie rate against Lfrqa as ground-truth as a metric to gauge systems’ RAG-QA quality. Figure 4 illustrates the evaluation framework.
The choice of Lfrqa as the target to compare has been partly justified in Sec. 3 for 1) Completeness: the annotation process encourages the inclusion of as much relevant information as possible. 2) Coherence: its answers are written more coherently and naturally than Robust-QA as references for LLM generations. Complete and coherent answers can be considered as a comprehensive summary of all relevant information in the entire corpus. This allows us to evaluate generated answers against Lfrqa answers only, which is much more informative and concise than using retrieved passages, potentially with a large amount of noise.
We implement human and model-based evaluations with the same instructions and report their correlations. We will show results in Sec. 6.1 that further justify using Lfrqa as evaluation targets.
Figure 4: Rag-qa Arena framework. Green blocks are LLM’s inputs to generate answers. Orange blocks are LLM and LFRQA’s answers presented to both human and LLM judges to determine pairwise preferences.
4.1 Human Evaluation
We present a query and a pair of answers (one from Lfrqa and one from an LLM), to human annotators. We instruct them to rate their preferences based on three aspects. 1. Helpfulness:  information that is helpful/relevant to answer the query. Touvron et al. (2023); Bai et al. (2022). 2. Truthfulness:  information that is correct to answer the query. By our definition, truthful information should also be helpful information Stephanie Lin (2021); Aisha Khatun (2024). 3. Completeness:  include as much truthful and helpful information as possible. We further instruct annotators to use Truthfulness (being both truthful and helpful) as the primary criterion since it is stricter than Helpfulness. Helpfulness is used when a decision cannot be made by Truthfulness alone. More details including the definition of aspects, rating categories, and step-by-step guidelines can be found in Appendix A.6 and Fig. 6-7 (annotation interface).
4.2 Model-Based Evaluation
Since human evaluation is too costly, we adopt model-based evaluators for scalable evaluation of LLMs on the entire Lfrqa test set.
As for the evaluation approach, we provide LLM-based evaluators with a query and a pair of answers (including one from Lfrqa). Similar to human evaluation, we prompt LLMs to rate their preferences based on the same three aspects above. We only modify the human instruction slightly to be compatible with LLM readable input text, but the majority of the prompt, especially the input data, and rubric, stay the same (Appendix Table 13-15).
For both human and model-based evaluations, we allow “tie" (no preference) as an option. For human evaluation, we take the majority votes from 3 annotators to mitigate biases. If there is no majority vote, we default the label to “tie.”
5 Experimental Setup
In this section, we discuss our retriever, LLMs experimented for both answer generation and pairwise evaluations and their prompts in more detail.
Retrieval setting.  We employ ColBERTv2 Santhanam et al. (2022) as our passage retriever, considering its superior performance on the underlying corpus for both Robustqa and Lfrqa as shown in Han et al. (2023). We follow the same retrieval setting and split passages into text chunks with 100 consecutive words. We use the top 5 retrieved passages for our main results in Table 3 and experiment with the top 10 passages for further analysis.
Answer generation.  We consider LLMs ranked top 25222Ranking at the time of paper writing. in the Chatbot Arena Chiang et al. (2024) and their smaller version to show the impact of model sizes. Due to resource and legal constraints, for proprietary LLMs, we only use OpenAI models: a) GPT-4-turbo (2024-04-09), b) GPT-4o and c) GPT-4-0125-preview). For public models, we experiment with 1) Mixtral-8x22B-Instruct and Mixtral-8x7B-Instruct Jiang et al. (2024); 2) Llama-3-70b-Instruct and Llama-3-8b-Instruct MetaAI (2024); 3) Command R+ and Command R Gomez (2024); 4) Qwen1.5-110b-chat and Qwen1.5-32b-chat Bai et al. (2023). Answer generation prompt can be found in Appendix Table 11.
Pairwise evaluation.  For LLM-based evaluators, we focus on a few larger models with strong context understanding capability, such as GPT-4-turbo, GPT-4-0125-preview, GPT-4o, Mixtral-8x22B-Instruct and Llama-3-70b-Instruct. Appendix Table 13-15 show the details of the pairwise evaluation prompts, including instruction, example prompt template, and in-context-learning examples. We shuffle the order of the answer pairs so that both human and model judges are not biased by the position of an answer. We select the LLM with the highest correlation with human judgments as the evaluator (Appendix Table 8). We use OpenAI API to run GPT-4 models. We download public models from HuggingFace Hub HuggingFace (2024) and run them on up to 8 Nvidia A100 GPUs with PyTorch (1.13.0) and Transformers (4.41.0) whose tokenizer.apply_chat_template() function can help adapt the generic prompts to different LLMs’ input formats.
We follow OpenAI’s recommendation [^3] https://platform.openai.com/docs/guides/prompt-engineering to design prompts with chain-of-thoughts (CoT) Wei et al. (2022), in-context learning Dong et al. (2023), and HTML tags as delimiters. We remove the thinking process in model outputs as final answers.
Overall [BI] [FI] [LI] [RE] [TE] [SC] [WR]
Compared Models W W+T W W+T W W+T W W+T W W+T W W+T W W+T W W+T
GPT-4o* #1 36.9 41.0 52.9 59.3 38.4 42.3 25.1 27.9 40.4 46.4 35.6 38.8 42.8 47.6 28.4 31.1
GPT-4-turbo #2 34.4 39.1 36.0 43.9 40.6 45.1 23.2 26.1 36.7 44.1 36.6 39.6 42.6 47.9 26.2 29.6
GPT-4-0125-preview #6 28.9 33.7 31.4 40.1 36.8 40.8 18.1 21.3 31.5 38.6 30.4 34.0 34.7 40.5 19.2 22.3
Mixtral-8x22b #3 34.5 38.8 37.0 46.0 44.1 47.6 21.3 24.4 34.4 41.0 33.9 36.8 45.0 49.5 25.9 28.1
Mixtral-8x7b #7 27.5 31.0 31.9 39.1 35.3 38.5 15.9 18.4 24.8 29.5 30.3 32.1 33.9 38.0 20.0 21.9
Llama-3-70b #8 21.7 25.2 30.3 37.2 24.6 27.7 12.9 15.1 22.3 27.3 22.4 24.4 25.6 30.0 15.5 18.2
Llama-3-8b #10 20.4 23.5 34.7 39.6 24.0 27.0 11.2 13.2 19.4 24.7 20.5 22.5 22.3 26.1 12.5 14.4
Command R+ #9 21.1 25.8 26.0 33.5 25.8 30.3 13.5 16.4 22.6 30.0 22.4 25.4 24.9 31.2 13.6 16.0
Command R #11 11.1 15.2 18.6 26.1 13.0 17.1 05.2 07.4 10.4 17.0 10.3 12.3 14.9 20.2 07.3 09.4
Qwen1.5-110b-chat #4 33.4 37.8 36.2 44.0 42.6 46.9 22.3 25.1 34.1 40.7 34.8 37.5 40.8 46.1 22.5 25.2
Qwen1.5-32b-chat #5 32.8 37.1 34.9 42.8 43.2 47.3 20.7 23.7 32.3 38.3 34.0 37.1 40.8 44.8 22.6 25.2
Table 3: Evaluation results on Lfrqa test set. W and W+T indicate win and win+tie rate against Lfrqa’s answers. LLM’s answers are generated based on the top 5 passages. bold and underline indicate the best and runner-up results. * means using the answer generation prompt w/o CoT. #n indicates the Elo ranking in Appendix Table 5.
Answer Pearson Cohen’s
Pairs Human MBE Corr. Kappa
Robustqa 06.1 01.0
(1) Lfrqa 83.9 95.6 0.82*** 0.71
Tie 10.0 03.4
GPT-4 78.1 77.1
(2) Robustqa 13.7 19.9 0.60*** 0.54
Tie 08.1 03.0
GPT-4 29.9 32.0
(3) Lfrqa 59.1 63.1 0.54*** 0.44
Tie 11.0 04.9
Mixtral 31.7 36.9
(4) Lfrqa 54.4 59.7 0.54*** 0.43
Tie 13.9 04.4
Llama-3 24.0 24.7
(5) Lfrqa 65.0 71.3 0.52*** 0.45
Tie 11.0 04.0
Table 4: Pairwise comparisons between Robustqa, Lfrqa, and LLMs. GPT-4: GPT-4-0125-preview; Mixtral: Mixtral-8x22b-Instruct; Llama-3: Llama-3-70b-Instruct. Answer generated based on top 5 passages using ColBERT-v2. MBE stands for the model-based evaluator. All numbers are % except for Pearson Corr. and Cohen’s Kappa. *** indicates strong correlation with p-values $\ll 0.001$.
6 Results and Analysis
Leveraging the evaluation framework described in Sec. 4, we first show that Lfrqa’s ground truth answers are dominantly preferred as answers than Robustqa. Then, we use the same evaluation framework to establish a new leaderboard, Rag-qa Arena, aiming to reliably measure RAG-QA systems’ performances across diverse domains.
(A) Lfrqa only (B) Lfrqa  + (C) Lfrqa  +
pairs 700 complete pairs 1400 complete pairs
Rag-qa Arena Ranking Rating 95% CI Votes Rating 95% CI Votes Rating 95% CI Votes
Lfrqa 1144 +1/-1 176.7K 1145 +1/-1 176.7K 1146 +1/-1 176.7K
GPT-4o 1066 +5/-5 16.1K 1081 +4/-4 23.1K 1085 +3/-3 30.1K
GPT-4-turbo 1050 +5/-4 16.1K 1058 +4/-3 23.1K 1065 +3/-2 30.1K
Mixtral-8x22b 1049 +4/-4 16.1K 1059 +3/-3 23.1K 1063 +3/-3 30.1K
Qwen1.5-110b-chat 1041 +4/-4 16.1K 1047 +4/-3 23.1K 1052 +3/-3 30.1K
Qwen1.5-32b-chat 1036 +6/-4 16.1K 1034 +4/-3 23.1K 1037 +3/-3 30.1K
GPT-4-0125-preview 1008 +6/-5 16.1K 1005 +4/-4 23.1K 1008 +3/-3 30.1K
Mixtral-8x7b 991 +4/-4 16.1K 991 +3/-4 23.1K 987 +3/-3 30.1K
Llama-3-70b 939 +4/-5 16.1K 931 +4/-4 23.1K 930 +2/-3 30.1K
Command R+ 938 +5/-5 16.1K 931 +3/-4 23.1K 924 +3/-3 30.1K
Llama-3-8b 924 +6/-6 16.1K 910 +4/-4 23.1K 903 +4/-3 30.1K
Command R 816 +8/-6 16.1K 802 +5/-5 23.1K 796 +4/-4 30.1K
Table 5: Elo rating including “couldn’t find answer” responses. “Lfrqa only” indicates the pairs that always include an Lfrqa answer. “Lfrqa  + N complete pairs” means we subsample N additional pairs evenly across 7 domains and conduct comparison for all pairs of models. We have 11 models, so the total new pairs are ${\frac{N\times 11\times 10}{2}}$.
6.1 LFRQA v.s. RobustQA
[…]
Table 4 shows that when compared directly in Row (1), Lfrqa dominates Robustqa. When comparing with GPT-4 in Row (2)-(3), Lfrqa significantly out-performs GPT-4, but Robustqa significantly under-performs. These results show strong evidence that Lfrqa’s answers can serve as better ground-truth than Robustqa.
6.2 Quality of Model-based Evaluator
To build Rag-qa Arena on the Lfrqa test set, we need a scalable evaluation method to benchmark various LLMs. We rely on model-based evaluation to achieve this goal. Before showing the final dashboard results, we check the quality of our selected evaluator (GPT-4-0125-preview) in Table 4.
To alleviate model bias, we use three LLMs’ answers as benchmark data, and the query set is the same 700 subsample above. Row (3)-(5) use GPT-4-0125-preview, Mixtral-8x22b-Instruct and Llama-3-70b-Instruct, respectively. All answers are generated based on the top 5 passages.
We observe that LLM evaluators’ numbers align well with the average human scores, except that LLMs tend to predict less “tie.” Most importantly, all Pearson Correlation Freedman et al. (2007) are above 0.52 (with p-values $\ll 0.01$), and all Cohen’s Kappa Cohen (1960) are above 0.43, both showing strong agreement between model and human judgments. In Appendix Table 8, we show correlation numbers using alternative LLMs as evaluators, but none of them works better than a single GPT-4-0125-preview model, which we select as our best quality evaluator for Rag-qa Arena.
6.3 RAG-QA Arena
Finally, we show Rag-qa Arena’s benchmark results. In Table 3 we report each model’s win and win+tie rate against Lfrqa.
Dashboard leaders.  GPT-4o leads the dashboard, with GPT-4-turbo and Mixtral-8x22b-Instruct as close runners-up. GPT-4o performs the best for [BI], [LI], [RE] and [WR] domains, Mixtral-8x22b-Instruct leads in [FI] and [SC], and GPT-4-turbo champions in [TE].
Impact of “no answer found.”  In RAG-QA, we rely on a passage retriever to provide context, which could be irrelevant. Our prompt (Appendix Table 11) asks an LLM to refrain from answering if it “couldn’t find an answer.” When we use this answer generation prompt with CoT (the last two lines in the prompt), GPT-4o produces 48.3% “I couldn’t find an answer” responses (Appendix Table 9). We randomly sample 20 such examples, and surprisingly found that in 16 cases, GPT-4o puts an answer in its <thinking> process, but continues to generate “I couldn’t find an answer.” Fig. 8-9 show four such examples in comparison with other LLMs’ answers with the same prompt, and GPT-4o’s new answers without CoT. As the answer generation prompt with CoT only fails for GPT-4o, we remove CoT for GPT-4o, which improves its answer format and reduces the “no-answer” ratio to the level similar to other competitive models.
These results raise a research question about the impact of prompt engineering. We emphasize that the goal of Rag-qa Arena is to propose a reliable evaluation framework, not to conduct extensive prompt engineering or model training to pursue the best RAG-QA system. We provide a dev set of Lfrqa, which can be leveraged in future research for model development purposes.
Elo rating.  Table 3 shows dashboard results of win and win+tie ratio against Lfrqa. We can further convert these pairwise comparisons into Elo ratings similar to Chatbot Arena Chiang et al. (2024). Table 5 reports our Elo ranking. The leftmost column shows Rag-qa Arena’s ranking based on win ratio: Lfrqa is ranked on the top followed by GPT-4o, GPT-4-turbo, etc. Column (A) uses the same data in Table 3, but here we only have comparisons between Lfrqa and other LLMs’ responses, i.e. there are no direct comparisons between LLMs’ responses. Thus, the total number of votes for are 176.7K for Lfrqa, and 16.1K for other LLMs. The new ranking based on Elo rating aligns with Rag-qa Arena, but the 95% confidence interval (CI) is not yet able to separate all different model pairs.
In Column (B), we add pairwise comparisons for all unique model pairs on 700 randomly sampled queries across seven domains (100 for each domain). This increases the votes for all LLMs compared to 23.1K. We continue to add 700 more queries in Column (C), which further increases the votes for all LLMs to 30.1K. We rank these pairs with the same LLM evaluator as in the main result table, and found that the ranking in (C) based on Elo rating aligns perfectly with Rag-qa Arena, and as we increase the pairs, the 95% CI can finally separate different models.
In general, the total added preference pairs are $\frac{N\times K\times(K-1)}{2}$, where $K$ is the number of models in Table 3, and $N$=700 and 1400 for Column (B) and (C), respectively. With 43.6% increase of total pairs (and thus the compute), the final ranking is identical with Rag-qa Arena based on win ratio, and only differs only slightly with the win+tie ranking in Table 3 by flipping the order of llama-3-70b and Command R+. These results present additional evidence that our approach of using Lfrqa only for pairwise comparisons is reliable. Furthermore, it reduces the computational costs from $\mathcal{O}(K^{2})$ to $\mathcal{O}(K)$ as we now only need to compare each LLM response once with the ground-truth in Lfrqa.
Impact of the number of passages.  In Appendix Table 10, we compare the top 3 LLMs by increasing the number of retrieved passages from 5 to 10. Doubling the number of passages (with extra costs) increases RAG-QA performances significantly. We also find that both GPT-4 models’ improvements are greater than Mixtral-8x22b-Instruct, showing their superior capability to understand long context and identify useful information from noise. The best win rate of GPT-4o against Lfrqa is 41.3%, which is 13.7% points lower than Lfrqa answers’ win rate against GPT-4o. This result shows that Lfrqa’s answer quality is difficult to surpass, further justifying using it as an evaluation target.
Impact of model sizes.  For the non-GPT LLM family, more parameters lead to better performances, but a larger increase in model sizes does not always indicate greater performance gains in our study. For example, the two Qwen1.5 models have the second-largest difference, but the lift from the smaller to the larger model is marginal. We leave more rigorous investigations to future research.
7 Discussion
7.1 LLM as Annotators
Citation
Completeness Accuracy Helpfulness
Lfrqa 90.8 88.9 48.1
GPT-4 75.3 65.5 35.2
Table 6: Comparison between human and GPT-4 annotations. All numbers are in %.
Using large language models to provide annotations has been explored in previous works Tan et al. (2024). It could provide a more scalable solution than human annotations but can suffer from hallucination and accuracy issues that require human validations Huang et al. (2023). We also experimented with LLM as annotators before we start human annotations. We subsample 100 queries from Lfrqa and prompt GPT-4-0125-preview to follow the similar procedure in Sec. 3.1 to combine answers (Appendix Table 12). Then we request our data linguists to compare Lfrqa and GPT-4 annotations based on 1) Completeness: whether all Robustqa answers are integrated into the final answers; 2) Citation Accuracy: whether citations in answers pointing to the right documents; 3) Helpfulness: defined the same as in Sec. 4. Table 6 shows Lfrqa out-performs GPT-4 annotations by 15.5%, 23.4% and 12.9% for the three dimensions, respectively, suggesting human annotations are both valuable and necessary for our task.
7.2 Alternative Evaluation Approaches
Using retrieved passages.  Rag-qa Arena leverages only Lfrqa’s annotations as ground-truth to directly evaluate LLM responses, and we explain this design choice in Sec. 4 that Lfrqa consists of complete and coherent answers that can be viewed as high-quality summary of all available answers in the entire corpus. This enables us to not show retrieved passages as they 1) increase the input length and thus the latency of an evaluator; and 2) they could contain incorrect information due to retrieval error, which mislead evaluators.
Using Lfrqa as references.  We can also use Lfrqa’s annotations as references when constructing the prompt for pairwise evaluation. That is, we can potentially compare a pair of LLMs’ responses by comparing them both against the references in a single trial. However, this approach would still require the similar $\mathcal{O}(K^{2})$ pairs as in the Elo rating, which is not as efficient as our proposed Rag-qa Arena framework.
For these reasons, we do not adopt the above two evaluation approaches. It is conceivable that prompt engineering, in-context example selections and even task specific evaluator training could further enhance alignments with human judges. We leave them for future research efforts.
8 Related Work
RAG-QA  has been widely studied. Prior datasets are limited in the evaluation as their corpus relies heavily on Wikipedia and the answers are mostly short and extractive Rajpurkar et al. (2016); Kwiatkowski et al. (2019); Amouyal et al. (2019). Robustqa and MultiHop-RAG Tang and Yang (2024) address the single domain issue, but still adopt short, extractive answers, which is not as suitable as Lfrqa to evaluate modern LLMs that generate long-form answers.
Longform QA datasets  have been proposed in prior work. ELI5 Fan et al. (2019) and LongFact Wei et al. (2024) contain answers that are either not annotated directly on the corpus, and or not created by humans. Krishna et al. (2021) also points out that ELI5’s small validation set has significant leakage from its train set. ASQA Stelmakh et al. (2022) is the most similar data to our work, but its corpus is in the single Wikipedia domain. Lfrqa is by far the RAG-QA dataset with the most comprehensive long-form answers.
Pairwise preference  is now a standard way to evaluate LLMs. It allows direct comparison between two responses Chiang et al. (2024); Lin et al. (2024). Rag-qa Arena is unique by always including a high-quality human annotated LFRQA answer, thereby making the evaluation more trustworthy.
9 Conclusion
We create Lfrqa, the first multi-domain dataset with coherent long-form answers to reliably benchmark RAG-QA. We propose a reliable LLM-based evaluation framework, Rag-qa Arena, that enables direct comparisons between LLMs’ answers and Lfrqa, which we believe will facilitate the evaluation RAG-QA in the era of LLMs.
Limitations
We discuss some limitations of this work for future research efforts. Rag-qa Arena can potentially cover more models. We didn’t include some leading LLMs, such as Claude Anthropic (2024) and Gemini Google (2023) models, due to legal and resource constraints, but we plan to add them to the leaderboard in the future. Evaluation using GPT-4-0125-preview is not cheap. It costs on average 300 U.S. dollars per model on the full Lfrqa’s test set. We plan to subsample 10-20% of the queries for the final public leaderboard, which will be more cost-friendly for future users. Future research can also study training smaller but equally accurate models as evaluators. Finally, we mainly focus on different LLMs for RAG-QA in this work, but future research can study the impact of different retrievers or joint retrievers and LLM training using Rag-qa Arena.
Ethics Statement
The authors of this paper are committed to conducting research ethically. We are leveraging existing LLMs to generate answers for Lfrqa, which include many open-ended questions. LLM-generated answers could be incorrect or unfaithful, as retrievers could find irrelevant passages and LLM can hallucinate Huang et al. (2023). These are known issues in the AI research community, and that is the reason we created Lfrqa to better evaluate RAG-QA systems. The additional risks and potential harms are discussed in numerous previous works Bender et al. (2021); Weidinger et al. (2021). The authors strive to ensure that the research and its results do not cause harm.
Data used in this work have been collected from public sources and used in accordance with all applicable laws and regulations. We use contracted data professionals for Lfrqa annotations, and Appen platform [^4] https://www.appen.com/ for human pairwise preference annotations. In both cases, we ensure our hourly rate is higher than 15 U.S. dollars per local minimum wage standard. The intended usage of Lfrqa is compatible with the underlying data’s access conditions (Appendix A.2)
Appendix A Appendix
Dev Set Robustqa Lfrqa
Domain Source Label $|Q|$ $|D|$ $|P|$ $A/Q$ $W/A$ $A/Q$ $W/A$
Lifestyle LoTTE [LI] 2,151 268,893 597,729 5.9 5.9 1.0 102.2
Recreation LoTTE [RE] 2,325 263,025 731,124 6.3 7.2 1.0 112.8
Technology LoTTE [TE] 2,223 1,000,000 1,707,346 5.4 8.9 1.0 83.3
Science LoTTE [SC] 2,137 343,642 854,756 4.8 6.4 1.0 100.6
Writing LoTTE [WR] 1,972 277,072 713,692 7.0 7.7 1.0 109.6
Table 7: Data summary for the dev set. Based on Robustqa, there is no dev split for BioASQ and FiQA data.
A.1 Dev Set of LFRQA
We do not use dev set to fit the scope of the paper, but we provide an additional 10K queries for future model developments, including prompting engineering and training. Details are shown in Table 7. All data are collected using the same process and quality control described in Sec. 3.
A.2 Data License
Lfrqa is created based on the following datasets. We make sure to follow the data distribution license for our usage of the data.
  • FiQA: no license provided, but all data information can be found one the official website [^5] https://sites.google.com/view/fiqa/home.
  • LoTTE: MIT license [^6] https://github.com/stanford-futuredata/ColBERT.
  • BioQAS: CC BY 2.5 license [^7] http://participants-area.bioasq.org/datasets/.
  • RobustQA: Apache-2.0 license [^8] https://github.com/awslabs/robustqa-acl23.
A.3 Annotation Interface
Our annotation interface can be found in Fig. 5. Annotators use this UI to write long-form answers. Blue highlights are the original answers from Robustqa.
A.4 Answer Generation Prompts
Table 11 shows our answer generation prompt. The last two lines “First, think step-by-step…” are what we refer to as the CoT prompt, which we remove for GPT-4o.
A.5 Evaluation Interface
We collect human pairwise preference data to benchmark our LLM-based evaluators. Fig. 6-7 show the details.
A.6 Human Evaluation Instructions
Helpfulness:  Information that is helpful/relevant to answer the query. An ideal answer consists of only information that is helpful/relevant to answer the query Touvron et al. (2023); Bai et al. (2022).
Truthfulness:  Information that is correct to answer the query. By our definition, truthful information should also be helpful information. If it is difficult to determine the truthfulness of some information, we consider it untruthful. Sometimes, this is due to not enough context provided in the answer. Another source of untruthfulness is when conflicting information is presented, and the answer does not coherently reconcile them Stephanie Lin (2021); Aisha Khatun (2024).
Completeness:  include as many helpful and truthful information.
Here are the details of our instructions.
1.  If one answer has all truthful information while the other has some untruthful information, prefer the all-truthful one. 2.  If both have some untruthful information, prefer the one with less untruthful information. 3.  If both have all truthful information, prefer the one with more truthful or helpful information. 4.  If two answers look equally good, or it is too hard to differentiate, choose “Not sure.”
As the annotation UI shows, the actual ratings are “Better,” “Slightly Better,” “Tie,” “Slightly Worse” and “Worse”. We merge “Better” and “Slightly Better,” and “Slightly Worse” and “Worse” when computing correlation with model-based evaluators.
A.7 Evaluation Model Selection
Table 8 shows Pearson Corr. between human and LLM-based evaluators. We find using GPT-4-0125-preview alone achieves the best outcome, and thus, use it for Rag-qa Arena evaluations.
Lfrqa v.s. Lfrqa v.s. Lfrqa v.s.
Evaluator Models GPT-4 Mixtral Llama-3
Llama-3-70b 0.52 0.51 0.48
Mixtral-8x22b 0.53 0.49 0.50
GPT-4-turbo 0.47 0.49 0.48
GPT-4o 0.43 0.46 0.45
GPT-4-0125-preview 0.54 0.54 0.52
Ensemble 0.53 0.53 0.52
Table 8: Correlation between human judges and leading LLMs as evaluators. Column titles are answer pairs. GPT-4: GPT-4-0125-preview; Mixtral: Mixtral-8x22b-Instruct-v0.1; Llama-3: Llama-3-70b-Instruct. The ensemble takes the majority vote out of the best three models.
A.8 No Answer Ratio
Table 9 shows the ratio of “I couldn’t find an answer" in each LLM’s answers.
A.9 Pairwise Evaluation Prompts
Table 13-15 show the details of the pairwise evaluation prompts including instruction, example prompt template, and in-context-learning examples.
A.10 Error Analysis
Fig. 8-9 show examples where the original GPT-4o puts an answer in its <thinking>, but generate “I couldn’t find an answer." It is fixed by removing CoT prompt. We also compare them with Mixtral-8x22b, GPT-4-turbo and Lfrqa’s answers.
Overall [BI] [FI] [LI] [RE] [TE] [SC] [WR]
GPT-4-turbo 14.1 10.9 19.8 12.2 15.9 11.2 12.3 12.1
GPT-4o 48.3 32.4 55.2 50.3 52.8 43.4 49.4 48.4
GPT-4o * 16.9 10.4 26.8 13.9 19.8 11.2 14.7 13.9
GPT-4-0125-preview 15.8 11.0 21.9 13.5 19.1 12.6 13.5 14.1
Mixtral-8x22b 9.9 10.1 13.6 9.1 13.7 7.1 6.6 6.0
Mixtral-8x7b 18.5 15.4 24.6 17.8 23.7 13.3 15.3 15.0
Llama-3-70b 25.7 17.7 35.9 24.4 29.2 20.5 23.3 21.2
Llama-3-8b 12.6 10.0 15.9 11.5 16.2 8.4 12.9 11.0
Command R+ 14.4 10.8 21.7 12.5 16.1 10.1 14.7 10.5
Command R 6.5 5.0 10.8 6.1 6.8 3.4 4.6 5.0
Qwen1.5-110b-chat 11.5 10.8 15.9 10.7 13.8 6.8 9.0 10.0
Qwen1.5-32b-chat 8.9 7.4 13.6 7.5 11.8 5.0 7.0 6.5
Table 9: No answer ratio for the entire Lfrqa’s test set. * means using answer generation prompt without CoT.
Overall [BI] [FI] [LI] [RE] [TE] [SC] [WR]
W W+T W W+T W W+T W W+T W W+T W W+T W W+T W W+T
GPT-4o* w/ 5 psgs 36.9 41.0 52.9 59.3 38.4 42.3 25.1 27.9 40.4 46.4 35.6 38.8 42.8 47.6 28.4 31.1
GPT-4o* w/ 10 psgs 41.3 45.0 59.1 64.4 45.7 49.4 27.4 30.0 43.7 49.9 39.6 42.2 46.0 50.0 30.7 33.1
GPT-4-turbo w/ 5 psgs 34.4 39.1 36.0 43.9 40.6 45.1 23.2 26.1 36.7 44.1 36.6 39.6 42.6 47.9 26.2 29.6
GPT-4-turbo w/ 10 psgs 40.6 45.4 40.8 49.5 49.3 53.7 27.5 30.7 41.9 49.5 44.2 47.8 48.3 53.1 31.6 34.4
Mixtral-8x22b w/ 5 psgs 34.5 38.8 37.0 46.0 44.1 47.6 21.3 24.4 34.4 41.0 33.9 36.8 45.0 49.5 25.9 28.1
Mixtral-8x22b w/ 10 psgs 38.1 42.5 41.8 50.0 47.4 51.3 24.9 28.0 39.1 46.0 38.9 41.1 46.7 51.5 28.1 30.7
Table 10: Impact of the number of passages. Evaluation results on the entire Lfrqa test set based on top 5 or 10 passages. W and W+T indicate win and win+tie rate against Lfrqa. bold and underline indicate the best and runner-up results. * means using the answer generation prompt w/o CoT.
Answer Generation Prompt
Based on the passages, provide a helpful answer to the query. Your answer must be faithful to the content in the passages.
Do not use your own knowledge to answer the query. If you couldn’t find any helpful information in the passages,
respond "I couldn’t find an answer."
Passages are inside <passage></passage> tags. Query is in the <query></query> tags.
{x.passages}
<query>
{x.question}
</query>
First, think step by step, and put your thinking in <thinking> tags. Your thinking must be shorter than 50 words.
Then, provide your answer.
Table 11: Prompt for answer generation. {x.*} indicates a component input that can be replaced by actual data. We modify this prompt slightly to be compatible with different LLMs’ input formats, but the majority of the prompt, particularly instructions, remain the same. * The last section starting with “First, think step by step…” is what we refer to as CoT prompt. We remove it for GPT-4o only.
Figure 5: Annotation Interface
Annotation Generation Prompt
Provide a response around 100 words to the query in the <query></query> tags based on the passages. Passages are
inside <passage></passage> tags. The response must incorporate all candidate answers in the <ans></ans>, and
you are allowed to rephrase these answers in order to make your final response natural. The response should not
include any information outside passages.
You should cite the passage number (indices) in the format of [1], [2], [3, 4], etc. at the end of each sentence.
{x.passages}
<query>
{x.question}
</query>
Table 12: Prompt for GPT-4 annotations. {x.*} indicates a component input that can be replaced by actual data.
Figure 6: Pairwise preference: instruction.
Figure 7: Pairwise preference: annotation interface.
Pairwise Evaluation: Instruction and Rubric
We will show you a query and a pair of answers to the query. You need to provide your preference over this pair
of answers.
First, try your best to determine whether the information in an answer can help truthfully answer the query. Then rate
your preference based on Helpfulness and Truthfulness.
- Helpfulness: information that is helpful/relevant to answer the query. An ideal answer consists of only information
                   that is helpful/relevant to answer the query.
- Truthfulness: information that you believe is correct to answer the query. By our definition, truthful information
                    should be helpful information. If you find it difficult to determine the truthfulness of some information,
                    consider it untruthful. Often time, this is due to not enough context provided in the answer. Another
                    source of untruthfulness is when conflicting information presented, and the answer does not reconcile
                    them in a coherent way.
<rubric>
Here is how you judge (in the order of importance),
- If one answer has all truthful information while the other has some untruthful information, prefer the all truthful one.
- If both have some untruthful information, prefer the one with less untruthful information.
- If both have all truthful information, prefer the one with more truthful or helpful information.
- If two answers look equally good, or it is too hard to judge using the 3 cases above, then you are our "not sure" which one
is better.
</rubric>
Table 13: Instruction and rubrics for pairwise evaluation. We use this template across all LLM evaluators. This can be considered as the "system" instruction for GPT-4 and LLama-3 models.
Pairwise Evaluation: Example Template
Query is in the <query></query> tags. Answer 1 is in <answer 1></answer 1>, and Answer 2 is in
<answer 2></answer 2>.
<query>
x.question
</query>
<answer 1>
x.response1
</answer 1>
<answer 2>
x.response2
</answer 2>
Review the rubric in <rubric> tags,
- if you prefer <answer 1>, output 1.
- if you prefer <answer 2>, output 2.
- if you are not sure, output 0.
First, think step by step, put your thinking in <thinking></thinking> tags. Your thinking must be shorter than 50 words.
Then, provide your rating inside <rating></rating> tags. Remember your rating should be 0 if you are not sure, and your
rating must be either 0, 1, or 2.
Table 14: Example template for pairwise evaluation. This template is used for both ICL examples and the final test example.
Label 1
Query difference between publicly and publically.
Answer 1 Both ‘publicly’ and ‘publically’ bear no difference in meaning, as they are essentially alternative spellings
of the same concept. Publicly is more widely used, but the existence of ’publically’ in reputable sources like
the OED means it cannot be dismissed as simply incorrect. Some opinions hold that ’publicly’ is the older
irregular form, still preached by a lot of grammars, and ’publically,’ on the other hand, is the newer and
regular form.
Answer 2 There is no difference in meaning between ’publicly’ and ’publically’; they are alternative spellings of the
same word.
Reason Both <answer 1> and <answer 2> are truthful. However, <answer 1> provides more truthful information
as the context to compare the two terms. Therefore, <answer 1> is better.
Label 2
Query what did European/American historical cooks do with the egg whites?
Answer 1 Historical European and American cooks used egg whites for making egg white omelettes and egg white
pasta, as well as for stiffening clothing, similar to how starch is used today.
Answer 2 Egg whites have found their place in various non-culinary applications throughout history, such as
in the clarification of beer and wine, in the conservation of books through bookbinding and gilding, and in
makeup as an ancient form of nail polish.
They were also utilized historically as a stiffening agent for clothing, similar to how starch is used today.
The culinary landscape was not left untouched, with egg whites making their way into recipes for omelettes
and pastas as early as the 15th century.
Reason Both <answer 1> and <answer 2> provide several usages of egg whites for European/American cooks.
<answer 2> clearly provides more options with fully explained details. Therefore, <answer 2> is better.
Label 0
Query should utf-16 be considered harmful?
Answer 1 The question of whether UTF-16 should be considered harmful is subject to differing opinions. One
perspective suggests that UTF-16 is harmful due to a historical misconception about character encoding and
recommends UTF-8 as the superior choice for various text interfaces. Another viewpoint argues that
UTF-16 is not inherently harmful, emphasizing its utility in specific scenarios where it serves as a
compromise between simplicity and compactness. The choice between UTF-16 and other encodings like
UTF-8 depends on the specific requirements of the application, such as compatibility with ASCII or
the need to efficiently encode certain character sets.
Answer 2 UTF-16 should not be considered harmful.
However, contrasting views argue that UTF-16 should indeed be considered harmful.
Some argue that the very reason UTF-16 exists is because some time ago there used to be a misguided belief
that WideChar is going to be what UCS-4 now is.
Additionally, the harmfulness of UTF-16 is tied to issues with exercising code.
Reason Both <answer 1> and <answer 2> reconcile the two conflicting views with detailed explanation.
I am not sure which one is better.
Table 15: In-context examples for pairwise evaluation. Labels 1, 2, and 3 mean "answer 1 is better", "answer 2 is better" and "tie", respectively. "Reason" is a model’s chain-of-thought output.
Figure 8: GPT-4o’s original “I couldn’t find an answer” response compared with GPT-4-Turbo and Mixtral-8x22b, and its answer without CoT (*). Blue highlights in the passages indicate helpful information to answer the query. <answer> tags are added to help differentiate from <thinking>.
Figure 9: GPT-4o’s original “I couldn’t find an answer” response compared with GPT-4-Turbo and Mixtral-8x22b, and its answer without CoT (*). Blue highlights in the passages indicate helpful information to answer the query. <answer> tags are added to help differentiate from <thinking>.